Is (-2,4) a solution to:
3x-2y=-14
5x+y=6

Found 2 solutions by ikleyn, rothauserc:
Answer by ikleyn(52781)

.
Is (-2,4) a solution to:


3x - 2y = -14?     Yes.    3*(-2) - 2*4 = -6 - 8 = -14.
5x +  y =   6?     No.     5*(-2) + 1*4 = -10 + 4 = -6


Answer by rothauserc(4718)

3x - 2y = -14
5x + y = 6
*****************************************************
we are asked to check the solution (-2, 4), namely x = -2, y = 4
1) 3x - 2y = -14
3(-2) - 2(4) = -14
-6 -8 = -14
-14 = -14
2) 5x + y = 6
5(-2) + 4 = 6
-10 + 4 = 6
-6 NOT = 6
******************************************************
therefore (-2, 4) is not a solution